OpenStudy (anonymous):

What value of b makes the polynomial below a perfect square? x2 + bx + 16 b = A. 16 B. 8 C. 2 D. 4

OpenStudy (anonymous):

(x + 4)^2 = x2 + bx + 16 Solve for b.

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow (x + 4)^2 }\) This is the value, clearly, as (x + 4)^2 will be x^2 + 16 + 2*4*x

OpenStudy (anonymous):

x^2 + 8x + 16 = x^2 + bx + 16 So b is 8.

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow (a + b)^2 = a^2 + 2a + b^2 }\)